# Efficient HDR Reconstruction from Real-World Raw Images

### v5

@@ 3.2.3 RealRaw-HDR Dataset
Kalanatri [14] 74 $1490\times 989$ sRGB 4 &\&& 8 &\&& 16
Chen [4] 144 $4096\times 2168$ raw, sRGB 4 &\&& 8 &\&& 16
@@
3.2.4 Effectiveness of the Data Formation Pipeline
The proposed pipeline for generating HDR data is efficient and user-friendly, allowing easy generation of numerous high-quality data pairs. Although our ground truth is derived from a single image, it contains a wide range of information characteristics of HDR images. Firstly, the raw images $I_{1}$ and $I_{2}$ are captured by a high dynamic range camera and contain high dynamic range. On the other hand, there is a significant difference in the signal-to-noise ratio between HDR images and LDR images. The exposure-aligned long-exposure images differ from the short-exposure images only in the dark and overexposed regions, in addition to the noise difference. Therefore, $I_{l}$ and $I_{s}$ have all the characteristics of real-world long and short-exposure images, and $I_{gt}$ contains a wide range of informative features of real HDR images.
4 Methodology
4.1 Overview
HDR reconstruction plays a vital role in various applications, such as mobile photography, high-definition displays, and virtual reality, where lightweight and efficient algorithms are highly demanded due to resource limitations. Previous learning-based HDR methods [21, 20, 33] often rely on large and complex models, making them impractical for real-world scenarios. On the other hand, unlike GPU servers, existing optimized manipulations for mobile devices are quite limited, especially on computationally limited devices. Unsupported operators have to be processed on the CPU, which not only very low processing speed, but also introduces additional MACs. Therefore, we first design a neat UNet with mobile-friendly operations as the base model, then propose a re-parameterizable Topological Convolution Block to improve HDR performance. For lightweight design, we do not use the computationally demanding explicit alignment in our HDR network. To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (in Sec. 4.5) that enables training with unaligned pairs.
4.2 Base Model
To ensure high inference speed and cross-device deployment on commodity mobile devices, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a neat UNet consisting of the most basic operations as the base model. The overall architecture of the base model is shown in Fig. 3(a), maximizing the use of the dual-exposure HDR sensor imaging characteristics. We introduce two distinct sub-encoders based on the differences in long- and short-exposure image features: Encoder-S and Encoder-L. Encoder-S extracts features from the short-exposure image, serving as reference features. In parallel, Encoder-L extracts features from the long-exposure image, offering supplementary features.
@@ 4.2 Base Model
Figure 4: Visual comparison of state-of-the-art HDR reconstruction methods on our synthetic test dataset.
4.3 Topological Convolution Block
Although the plain base model is efficient, its HDR performance is less satisfactory compared to those complicated models, as shown in Tab. 5. We thus employ the re-parameterization technique to enrich the representation capability of the base model. The reparameterization has achieved promising results on other tasks [7, 3, 5, 44]. We design a flexible re-parameterizable module called the Topological Convolution Block (TCB)), which can more effectively extract edge and texture information for the HDR task. As shown in Fig. 3(b), the TCB consists of several fundamental units: (1) A standard $3\times 3$ convolution for a solid foundation. The standard convolution is denoted as:
@@ 4.3 Topological Convolution Block
(2) Extending and squeezing convolution to enhance feature expressiveness, the expanding and squeezing feature is extracted as:
@@
(3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges, i.e., using a predetermined convolution kernel to process the edge information. Denote by $D_{x}$ and $D_{y}$ the horizontal and vertical Sobel filters, $D_{lap}$ is the Laplacian filter.
@@
The output of the TCB in the combination of the five components:
$F_{TCB}=F_{n}+F_{es}+F_{edge}+F_{c}+F_{j}$ (14)
@@
4.4 Re-parameterization for Efficient Inference
To achieve an efficient HDR network that meets the stipulated design prerequisites of low computational complexity and streamlined hardware device deployment, we simplify the TCB reparameterization into a single $3\times 3$ convolution after training. Following previous works [6, 5, 44], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolution into a single $3\times 3$ convolution. Additionally, we combine the Sobel and Laplacian operators into a special $3\times 3$ convolution with a fixed convolution kernel. The $1\times 1$ convolution is achieved by padding the convolution kernel with zeros. As a result, TCB can be transformed into a $3\times 3$ convolution for efficient implementation during the inference stage, as shown in Fig. 3(b). By utilizing TCB, we achieve superior HDR results with improved efficiency.
@@ 4.4 Re-parameterization for Efficient Inference
Table 2: Performance comparison of different HDR models on our synthetic dataset. #Param and FLOPs represent the total number of network parameters and floating-point operations. The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images. Metrics with $\uparrow$ and $\downarrow$ denote higher better and lower better, respectively. The best and second-best performances are in bold and underlined, respectively. "-" indicates the result is not available.
Methods FLOPs #Param Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
DeepHDR[33] 2409.32G 15.26M 4.3 ms 43.3680 1.3767 43.5551 1.3844 43.7312 1.3679 42.8178 1.3779
NHDRRNet[37] 826.17G 40.26M 7.9 ms 33.0206 2.7308 33.0127 2.7277 33.0392 2.7203 33.0101 2.7443
UNet-SID[2] 640.89G 7.76M 3.1 ms 43.3892 1.3434 43.3314 1.3535 43.4551 1.3312 43.3811 1.3456
SGN[8] 712.66G 4.78M 3.3 ms 43.6094 1.3235 43.5074 1.3398 43.7078 1.3067 43.6131 1.3240
HDR-Transformer[21] 3698.28G 1.23M - 44.3895 1.3681 44.3311 1.3811 44.4140 1.3619 44.4235 1.3613
AHDRNet[34] 2848.29G 0.93M 23.6 ms 44.7985 1.2939 44.8548 1.2957 44.8343 1.2892 44.7064 1.2968
Ours 127.55G 0.82M 2.9 ms 44.8081 1.2886 44.7575 1.3000 44.8482 1.2812 44.8187 1.2842
4.5 Loss Functions
@@
Based on the above analysis, we devise the strategies for dual-exposure HDR fusion: Firstly, in scenarios involving motion or overexposure within the fused region, we prefer to select short-exposure image that contains more information; Secondly, when the SNR of the short frame is too low, we prefer to select the long-exposure image that contains more information; Thirdly, our strategy is inclined to address ghost artifacts with a higher priority than lower SNR when ghost artifacts and lower SNR concurrently exist.
@@
To achieve the best HDR reconstruction results, we employ the multi-scale structural similarity loss function loss guide model, which learns short-exposure image information for global motion. By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
@@
Bayer loss. We propose a color correction loss, named Bayer loss, to minimize color cast and artifacts. We average the two G channels of the output (RGGB pattern) and ground truth (RGGB pattern) respectively, and then concatenate the averaged G channel with the R and B channels to perform a naive transformation to the RGB color space, producing two RGB images: $\widetilde{I}^{out}_{rgb}$ and ${I}^{gt}_{rgb}$. Then, we impose the colorfulness loss between the processed output and ground truth by the cosine embedding loss.
@@
Figure 6: Visual comparisons with the state-of-the-art methods on HDR sensor raw dataset.
Table 3: Performance comparison of different HDR models on the actual HDR sensor raw dataset [4]. The best and second-best performances are in bold and underlined, respectively.
@@
5 Experiments
5.1 Experimental Setup
We perform a quantitative evaluation using the PSNR, SSIM, and CIE L*a*b* space  [^1] CIE L*a*b* is a color space specified by the International Commission on Illumination. It describes all the colors visible to the human eye and was created to serve as a device-independent model for reference. [43, 11, 19] (also known as $\Delta E$). $\Delta E$ can effectively assess chromaticity, contrast, and color accuracy variations within HDR images rather than exclusively concentrating on luminance differences (HDR-VDP-2). It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE L*a*b* color space.
$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (20)
where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE L*a*b* version of the predicted HDR image and ground truth, respectively.
Implementation details. We train our model using the Adam optimizer [15] with weight decay $1\times 10^{4}$, learning rate $10^{-4}$, and $\beta_{1}$ and $\beta_{2}$ values set to 0.9 and 0.999, respectively. The input patch size for the network is $256\times 256$, and the batch size is 32. Our model is implemented in PyTorch and trained with an NVIDIA RTX 3090 GPU.
5.2 Comparison with the Other Methods
We choose several representative low-level vision methods for comparisons, including four HDR methods based on sRGB images (AHDRNet [34], DeepHDR [33], NHDRRNet [37], HDR-Transformer[21]), as well as two methods for denoising raw images (SGN [8] and UNet-SID [2]). For fair comparisons, we re-train all the methods using the RealRaw-HDR dataset. Additionally, for AHDRNet, DeepHDR, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images. Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure raw pairs as inputs.
Evaluation on synthetic dataset. We first evaluated our method on a synthetic dataset generated using the raw HDR data formation pipeline. The quantitative comparison results are shown in Tab. 2. The results clearly show that our method outperforms previous methods in almost all metrics on the synthetic dataset. Notably, our lightweight-efficient RepUNet model has fewer parameters (0.82M) and remarkably low GFLOPs (127G FLOPs). This efficiency allows us to process two 4K Bayer raw images at only 2.9ms using an NVIDIA RTX 3090 GPU. In contrast, other models with comparable performance necessitate significantly longer processing times. RepUNet achieves comparable performance with AHDRNet with only $4.5\%$ of its computational complexity (127G vs. 2848G). Fig 4 visualizes that our method can effectively eliminate noise and ghosting artifacts in the reconstructed HDR. In comparison, DeepHDR [33], NHDRRNet [37], and SGN [8] exhibit numerous artifacts in the palm motion region. However, our proposed RepUNet can reconstruct HDR images without ghosting (see rows 2 in Fig. 4).
Figure 7: Visual comparison of state-of-the-art HDR reconstruction methods on FUJI raw dataset.
Table 4: Quantitative comparisons of different loss functions. AMSS-Loss represents the alignment-free and motion-aware short-exposure-first selection loss.
@@ 5.2 Comparison with the Other Methods
Table 5: Reparameterization ablation results. The FLOPs and run times are measured on the raw image with a 4K resolution.
Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]

### v6

@@ III-C RealRaw-HDR Dataset
Kalanatri [18] 74 $1490\times 989$ sRGB 4 &\& 8 &\& 16
Chen [5] 144 $4096\times 2168$ raw, sRGB 4 &\& 8 &\& 16
@@
III-D Effectiveness of the Data Formation Pipeline
The proposed pipeline for generating HDR data is efficient and user-friendly, allowing easy generation of numerous high-quality data pairs. Although our ground truth is derived from a single image, it contains a wide range of information characteristics of HDR images. Firstly, the raw images $I_{1}$ and $I_{2}$ are captured by an HDR camera. On the other hand, there is a significant difference in the signal-to-noise ratio between HDR images and SDR images. The exposure-aligned long-exposure images differ from the short-exposure images only in the dark and overexposed regions, in addition to the noise difference. Therefore, $I_{l}$ and $I_{s}$ have all the characteristics of real-world long and short-exposure images, and $I_{gt}$ contains a wide range of informative features of real HDR images.
IV Methodology
IV-A Overview
HDR reconstruction plays a vital role in various applications, such as mobile photography, high-definition displays, and virtual reality, where lightweight and efficient algorithms are highly demanded due to resource limitations. Previous learning-based HDR methods [25, 24, 44] often rely on large and complex models, making them impractical for real-world scenarios. On the other hand, unlike GPU servers, existing optimized manipulations for mobile devices are quite limited, especially on computationally limited devices. Unsupported operators have to be processed on the CPU, which not only very low processing speed but also introduces additional MACs. Therefore, we first design a neat UNet with mobile-friendly operations as the base model, then propose a reparameterizable Topological Convolution Block to improve HDR performance. For lightweight design, we do not use the computationally demanding explicit alignment in our HDR network. To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (in Sec. IV-E) that enables training with unaligned pairs.
IV-B Base Model
To pursue high inference speed on commodity mobile devices and the possibility of cross-device deployment, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a neat UNet consisting of the most basic operations as the base model. The overall architecture of the base model is shown in Fig. 3(a), maximizing the use of the dual-exposure HDR sensor imaging characteristics. We introduce two distinct sub-encoders based on the differences in long- and short-exposure image features: Encoder-S and Encoder-L. Encoder-S extracts features from the short-exposure image, serving as reference features. In parallel, Encoder-L extracts features from the long-exposure image, offering supplementary features.
@@ IV-B Base Model
IV-C Topological Convolution Block
Although the plain base model is efficient, its HDR performance is less satisfactory compared to the complicated models, as shown in Tab. VIII. We thus employ the reparameterization technique to enrich the representation capability of the base model. The reparameterization has achieved promising results on other tasks [8, 4, 6, 55]. We design a flexible reparameterizable module called the Topological Convolution Block (TCB), which can more effectively extract edge and texture information for the HDR task. As shown in Fig. 3(b), the TCB consists of several fundamental units:
(1) A standard $3\times 3$ convolution for a solid foundation. The standard convolution is denoted as:
@@ IV-C Topological Convolution Block
(2) Extending and squeezing the convolution to enhance feature expressiveness, the expanding and squeezing feature is extracted as:
@@
(3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges, i.e., using a predetermined convolution kernel to process the edge information. Denote by $D_{x}$ and $D_{y}$ the horizontal and vertical Sobel filters, and $D_{lap}$ is the Laplacian filter.
@@
The output of the TCB is the combination of the five components:
$F_{TCB}=F_{n}+F_{es}+F_{edge}+F_{c}+F_{j}.$ (14)
@@
IV-D Re-parameterization for Efficient Inference
To achieve an efficient HDR network that meets the stipulated design prerequisites of low computational complexity and streamlined hardware device deployment, we simplify the TCB reparameterization into a single $3\times 3$ convolution after training. Following previous works [7, 6, 55], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolutions into a single $3\times 3$ convolution. Additionally, we combine the Sobel and Laplacian operators into a special $3\times 3$ convolution with a fixed convolution kernel. The $1\times 1$ convolution is achieved by padding the convolution kernel with zeros. As a result, TCB can be transformed into a $3\times 3$ convolution for efficient implementation during the inference stage, as shown in Fig. 3(b). By utilizing TCB, we achieve superior HDR results with improved efficiency.
@@ IV-D Re-parameterization for Efficient Inference
IV-E Loss Functions
@@
Based on the above analysis, we devise the strategies for dual-exposure HDR fusion: Firstly, in scenarios involving motion or overexposure within the fused region, we prefer to select the short-exposure image that contains more information; Secondly, when the SNR of the short frame is too low, we prefer to select the long-exposure image that contains more information; Thirdly, our strategy is inclined to address ghost artifacts with a higher priority than lower SNR when ghost artifacts and lower SNR concurrently exist.
@@
To achieve the best HDR reconstruction results, we employ the multi-scale structural similarity loss function guide model, which learns short-exposure image information for global motion. By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
@@
Bayer loss. We propose a color correction loss, named Bayer loss, to minimize color cast and artifacts. We average the two G channels of the output (RGGB pattern) and ground truth (RGGB pattern) respectively, and then concatenate the averaged G channel with the R and B channels to perform a naive transformation to the RGB color space, producing two RGB images: $\widetilde{I}^{out}_{rgb}$ and ${I}^{gt}_{rgb}$. Then, we impose the colorfulness loss between the processed output and the ground truth by the cosine embedding loss.
@@
TABLE II: Performance comparison of different HDR models on synthetic dataset from RealRaw-HDR. #Param and FLOPs represent the total number of network parameters and floating-point operations. The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images. Metrics with $\uparrow$ and $\downarrow$ denote higher better and lower better, respectively. PSNR test in HDR Raw images and $\Delta E$ test in tone-mapped sRGB images. The best and second-best performances are in bold and underlined, respectively. ”-” indicates that inference is not possible due to memory limitations.
Methods GFLOPs #Param Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
DeepHDR[44] 2409.32 15.26M 4.3ms 43.3680 1.3767 43.5551 1.3844 43.7312 1.3679 42.8178 1.3779
NHDRRNet[49] 826.17 40.26M 7.9ms 33.0206 2.7308 33.0127 2.7277 33.0392 2.7203 33.0101 2.7443
UNet-SID[3] 640.89 7.76M 3.1ms 43.3892 1.3434 43.3314 1.3535 43.4551 1.3312 43.3811 1.3456
SGN[10] 712.66 4.78M 3.3ms 43.6094 1.3235 43.5074 1.3398 43.7078 1.3067 43.6131 1.3240
CEN-HDR[36] 162.35 0.19M 2.8ms 42.3021 1.4240 42.5312 1.4237 42.6531 1.4254 41.7219 1.4228
HDR-GAN[31] 629.09 1.32M 8.3ms 44.7641 1.2972 44.8954 1.2867 44.7743 1.3272 44.6427 1.2778
HDR-Transformer[25] 3698.28 1.23M - 44.3895 1.3681 44.3311 1.3811 44.4140 1.3619 44.4235 1.3613
AHDRNet[46] 2848.29 0.93M 23.6ms 44.7985 1.2939 44.8548 1.2957 44.8343 1.2892 44.7064 1.2968
SCTNet[37] 3025.80 0.97M - 44.8452 1.3048 44.9132 1.3231 44.8712 1.3026 44.7511 1.2889
Ours 127.55 0.82M 2.9ms 44.8081 1.2886 44.7575 1.3000 44.8482 1.2812 44.8187 1.2842
Figure 5: Visual comparison of state-of-the-art HDR reconstruction methods on our synthetic dataset from RealRaw-HDR.
V Experiments
V-A Experimental Setup
We perform a quantitative evaluation using the PSNR, SSIM, HDR-VDP3 [28], PU [2], ColorVideoVDP [29], and CIE L*a*b* space. The HDR-VDP3 [28] metric predicts the quality degradation concerning the reference image. We adopt the default settings provided in the HDR-VDP-3 processing framework, which simulate a typical HDR viewing environment. Metrics are computed directly on the HDR images without any tone mapping. All HDR images are evaluated under a display profile with peak luminance of 1000 cd/m², a contrast ratio of 1000:1, a gamma of 2.2, an ambient illumination of 100 lux, and a screen size of 30 inches at $3840\times 2160$ resolution. The viewing distance was set to 0.5 meters. To further validate the perceptual quality of the reconstructed HDR images, we incorporated the ColorVideoVDP [29] into our evaluation. We used the official implementation (v0.5.0) and applied it to HDR images using the configuration: ”ColorVideoVDP v0.5.0, 75.4 [pix/deg], Lpeak=1000, Lblack=0.5979 [cd/m²], (standard_4k).” CIE L*a*b* space  [^1] CIE L*a*b* is a color space specified by the International Commission on Illumination. describes all the colors visible to the human eye and was created to serve as a device-independent model for reference. [57, 14, 23] (also known as $\Delta E$). It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE L*a*b* color space.
$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (20)
where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE L*a*b* version of the predicted HDR image and ground truth, respectively. Note that HDR-VDP3, PU, and $\Delta E$ are all tested on sRGB images, which are obtained from HDR Raw images through the same image signal processor process.
TABLE III: Performance comparison of different HDR models on the actual HDR sensor raw dataset from Chen’s dataset [5]. PSNR and SSIM tests in HDR Raw images. The best and second-best performances are in bold and underlined, respectively.
@@ V-A Experimental Setup
TABLE IV: Performance comparison of different HDR models on the actual HDR sensor raw dataset from Chen’s dataset [5]. All HDR images are now evaluated under a 1000 $cd/m^{2}$ display profile, with metrics computed on the HDR image without tone mapping. The best and second-best performances are in bold and underlined, respectively.
Methods All-Exposure Ratio=4 Ratio=8 Ratio=16
@@
DeepHDR[44] 9.034 38.180 9.646 9.042 37.581 8.939 39.136 9.121 37.824
NHDRRNet[49] 8.771 30.743 9.276 8.687 30.756 8.739 30.946 8.885 30.567
UNet-SID[3] 9.089 38.083 9.701 9.087 37.613 8.986 38.969 9.194 37.665
SGN[10] 9.104 37.696 9.705 9.074 37.376 9.010 38.368 9.230 37.344
HDR-Transformer[25] 9.106 38.145 9.697 9.093 37.694 9.028 38.752 9.197 37.988
AHDRNet[46] 9.153 38.220 9.734 9.189 37.876 9.037 38.756 9.234 38.038
Ours 9.138 39.276 9.702 9.108 39.459 9.060 39.919 9.247 38.449
Figure 6: Visual comparisons with the state-of-the-art methods on the actual HDR sensor raw dataset from Chen’s dataset [5].
Implementation details. We train our model using the Adam optimizer [19] with weight decay $1\times 10^{4}$, learning rate $10^{-4}$, and $\beta_{1}$ and $\beta_{2}$ values set to 0.9 and 0.999, respectively. The input patch size for the network is $256\times 256$, and the batch size is 32. Our model is implemented in PyTorch and trained with an NVIDIA RTX 3090 GPU with Intel Xeon Platinum 8369B CPU (64 vCPUs@2.90GHz). The total training time is 60 hours.
V-B Comparison with the Other Methods
We choose several representative low-level vision methods for comparisons, including six HDR methods based on sRGB images (AHDRNet [46], DeepHDR [44], NHDRRNet [49], HDR-GAN [31], HDR-Transformer [25], CEN-HDR [36], and SCTNet [37]), as well as two methods for denoising raw images (SGN [10] and UNet-SID [3]). For fair comparisons, we retrain all the methods using the RealRaw-HDR dataset. Additionally, for AHDRNet, DeepHDR, HDR-GAN, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images. Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure raw pairs as inputs.
Evaluation on the synthetic dataset. We first evaluated our method on a synthetic dataset generated using the raw HDR data formation pipeline. The quantitative comparison results are shown in Tab. II. The results demonstrate that our method achieves performance comparable to the state-of-the-art (SOTA) method on the synthetic dataset. Notably, our lightweight and efficient RepUNet architecture significantly reduces both model size and computational complexity, with only 0.82M parameters and 127 GFLOPs. This high efficiency enables the processing of 4K Bayer raw images in just 2.9 ms on an NVIDIA RTX 3090 GPU, outperforming other models with similar performance that require substantially longer processing times. Compared to CEN-HDR, RepUNet improves PSNR by 2.5 dB while maintaining comparable computational costs. Furthermore, RepUNet achieves performance on par with AHDRNet, SCTNet, and HDR-GAN, but with dramatically lower computational complexity: it requires only 4.5% of the computational effort of AHDRNet (127 GFLOPs vs. 2848 GFLOPs), 4.2% of SCTNet (127 GFLOPs vs. 2025 GFLOPs), and 20.2% of HDR-GAN (127 GFLOPs vs. 629 GFLOPs). These results highlight the superior efficiency and practicality of our approach. Fig. 5 shows that our method can effectively eliminate noise and ghosting artifacts in the reconstructed HDR. In comparison, DeepHDR [44], NHDRRNet [49], and SGN [10] exhibit numerous artifacts in the palm motion region. However, our proposed RepUNet can reconstruct HDR images without ghosting (see row 2 in Fig. 5).
Figure 7: Visual comparison of state-of-the-art HDR reconstruction methods on the FUJI Raw dataset, captured with a Fujifilm GFX50S II camera containing real-world bracketed-exposure raw images.